OpenStudy (anonymous):

Which of the following is equivalent to |5x – 1| = 20?

OpenStudy (anonymous):

What are your choices?

OpenStudy (anonymous):

5x – 1 = 20 5x – 1 = 20 or 5x – 1 = -20 5x – 1 = -20 none of the above

OpenStudy (anonymous):

Okay, absolute value, or the | | marks, mean that there will be no negatives. This equation is saying that 5x - 1 = 20.

jimthompson5910 (jim_thompson5910):

if |x| = k then x = k or x = -k for any positive number k

OpenStudy (anonymous):

if you ever see anything that is like \[\left| x \right|\] it is always going to be a positive number